OpenStudy (mtalhahassan2):

a) Determine the equation of the inverse of the quadratic function(x) = x square -4x+3. b) List the domain and range of f(x) and its inverse. c) Sketch the graph of f(x) and its inverse.

OpenStudy (mtalhahassan2):

Please help me with this question!@

OpenStudy (campbell_st):

ok... so swap x and y \[x = y^2 - 4y + 3\] so looks like you should subtract 3 then complete the square... as a starting process

OpenStudy (mtalhahassan2):

bro why don't we have to put the equation in vertex foam

OpenStudy (campbell_st):

it's not necessary... it's a question with a variety of methods available to solve it.

OpenStudy (mtalhahassan2):

so what is the next step

OpenStudy (campbell_st):

ok... so subtract 3 and you get \[x - 3 = y^2 - 4y\] complete the square in y on the right side and add the value to both sides.

OpenStudy (campbell_st):

just post you're solution to the next step

OpenStudy (mtalhahassan2):

ok

OpenStudy (mtalhahassan2):

But how can we find the domain and range?

OpenStudy (campbell_st):

well start by finding the inverse...

OpenStudy (mtalhahassan2):

Please tell me about that ?

OpenStudy (campbell_st):

it's easier to see and understand the domain and range of the inverse, after you have found the inverse.

OpenStudy (campbell_st):

so what domain and range do you have for the original function?

OpenStudy (mtalhahassan2):

2,3

OpenStudy (campbell_st):

nope... the domain are the x values you can input into the function. the range are the output values.

OpenStudy (mtalhahassan2):

So it we the answer for inverse function: x - 3 = y^2 - 4y

OpenStudy (mtalhahassan2):

I am confuse

OpenStudy (mtalhahassan2):

What we do after this x - 3 = y^2 - 4y

OpenStudy (campbell_st):

nope that's the start you need to complete the square on the right hand side so its a factored perfect square.

OpenStudy (mtalhahassan2):

but who

OpenStudy (mtalhahassan2):

Can you please show me?

OpenStudy (mtalhahassan2):